yaquii:

Anonymous asked in Science & MathematicsPhysics · 5 years ago You stand on a bridge above a river and drop a rock into the water below from a height of 53 m. (Assume no air resistance.)? a. What is the acceleration of the rock as it falls? b. How long does it take to hit the water? c. How much more time would it take if instead of dropping the rock straight down, you threw it horizontally with a speed of 5 m/s? Explain. d. You throw another rock at an angle of 30° above the horizontal with a speed of 12 m/s. What are the x- and y-components of the rock's speed? e. How long does it take the rock to reach its maximum height?

Vocaloid:

a. the only force acting in the y-direction is **gravity** so think acceleration due to gravity - what's the constant and value for this? b. kinematics equation delta y = v0 * t + 1/2 at^2 since you're dropping it, rather than throwing it, there's no initial y velocity. so the equation reduces to delta y = 1/2 at^2 where delta y is the height (53 m), a is the acceleration due to gravity, and t is time. solve for time. c. think - does adding a horizontal force affect anything about our previous calculation? as a major hint look at what is/isn't in the equation. d. component form: x-component = v0*cos(theta), y-component = v0*sin(theta) - at (have to account for gravity). plug in and solve. e. there's a couple of ways you could solve it but the way I would do it is to use the kinematics equation in the y-direction, plug in vf (in the y direction) = 0 and solve for t

yaquii:

i didn't understand

Markg720:

Yeah my answer was kinda the answer but she took it down

Markg720:

Sorry brotha

Vocaloid:

may I ask what part you got stuck on?